12. A car travels a distance of 250 miles, 700 miles and 325 miles at the rate of 50 miles/hr, 35 miles/hr and 13 miles/hr respectively.Find the average speed of the car in miles/hr.
A. 30
B. 25.5
C. 42.5
D. 13
E. 50

Hello!
Can you please show me step by step how to solve the answer.
Thank you
Tania

Answer by mananth(16946)

A car travels a distance of 250 miles, 700 miles and 325 miles at the rate of 50 miles/hr, 35 miles/hr and 13 miles/hr respectively.Find the average speed of the car in miles/hr.
A. 30
B. 25.5
C. 42.5
D. 13
E. 50

250 miles at 50 mph
time taken = 250/50 = 5 hours------------ time = distance / rate
700 miles at 35 mph
time taken = 700/35= 20 hours
325 miles at 13 mph
time taken = 325/13 = 25 hours
total distance traveled at all three speeds = 250+700+325=1275 miles
Total time taken to travel 1275 miles = 50 hours

The average speed will be 1275 /50 = 25.5 mph